OpenStudy (anonymous):

What are the zeros of the function f(x) =3x^2-3x-6/3x-6

OpenStudy (anonymous):

you mean f(0)?

OpenStudy (anonymous):

no f(x)

OpenStudy (lgbasallote):

the zeroes would be simply when f(x) = 0 then solve for the value of x

OpenStudy (lgbasallote):

\[\frac{3x^2 - 3x-6}{3x - 6} = 0\] solve for x

OpenStudy (lgbasallote):

does that help? or do you need more help?

OpenStudy (anonymous):

i need more help idont know how to solve that

OpenStudy (lgbasallote):

start by factoring out the numerator and denominator. what do you get?

OpenStudy (anonymous):

3x^2=0 right

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

you cannot cancel 3x-6 that easily

OpenStudy (lgbasallote):

you factor it this way \[\frac{3(x^2 - x - 6)}{3(x-2)} = 0\] got it?

OpenStudy (lgbasallote):

uhh wait that should be 2

OpenStudy (lgbasallote):

\[\frac{3(x^2 - x - 2)}{3(x-2)} = 0\]

OpenStudy (lgbasallote):

now what can you cancel?

OpenStudy (anonymous):

x-2 and 3

OpenStudy (lgbasallote):

you cant cancel x-2

OpenStudy (lgbasallote):

when you have addition you cant cancel

OpenStudy (lgbasallote):

when you have x^2 - x - 2 and x - 2 you cant cancel x-2

OpenStudy (lgbasallote):

because there is x^2

OpenStudy (lgbasallote):

so 3 is the only thing that you can cancel \[\frac{x^2 - x - 2}{x-2} = 0\] now factor out the numerator

OpenStudy (anonymous):

idk im just trying to pass the class this is my last lesson so ill go with -1

OpenStudy (lgbasallote):

just factor out the numerator

OpenStudy (lgbasallote):

what is the factored form of x^2 - x -2

OpenStudy (lgbasallote):

you know how to factor right?